OpenStudy (anonymous):

If the product of two positive numbers is 1 and their difference is (5/6) what is the larger number?

OpenStudy (mertsj):

ab=1 a-b=5/6 Solve the system

OpenStudy (anonymous):

my answer comes out wrong when i try to solve the system

OpenStudy (anonymous):

does it need to be a quadratic?

OpenStudy (mertsj):

b=1/a from the first equation Substitute into the second equation. a-1/a=5/6 Multiply both sides by a a^2-1=5/6(a) Multiply both sides by 6 6a^2-6=5a Subtract 5a from both sides. 6a^2-5a-6=0 Factor and solve